# FERMILAB-PUB-25-0130-PPD

### v1

@@ Abstract
Understanding electron neutrino interactions is crucial for measurements of neutrino oscillations and searches for new physics in neutrino experiments. We present the first measurement of the flux-averaged and charged current single charged pion production cross section on argon using the MicroBooNE detector and data from the NuMI neutrino beam. The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon at a mean and energy of 730 MeV. Differential cross sections are also reported in electron energy, electron and pion angles, and electron-pion opening angle.
Introduction.— The next generation of accelerator neutrino oscillation experiments will seek to address multiple open questions in neutrino physics through precision measurement of electron neutrino appearance in muon neutrino beams. These include the presence and scale of charge-parity violation in the neutrino sector, the neutrino mass hierarchy, and the resolution of anomalies observed at short baselines [1]. Several of these experiments will make use of the liquid argon time projection chamber (LArTPC) detector technology [2, 3]. As such, in order to facilitate these measurements, precise understanding of interactions on argon is essential.
Cross section modeling of interactions is typically extrapolated from measurements. However, uncertainties on the interaction cross section ratio limit their constraining power [4, 5, 6, 7]. Understanding the nuclear effects unique to -argon interactions is crucial for precision neutrino oscillation measurements, rare-event searches, and benchmarking theoretical models used in neutrino experiments.
@@
Existing measurements of cross sections on argon are limited consisting of several inclusive measurements [8, 9, 10] and a measurement without final state pions [11]. The Deep Underground Neutrino Experiment (DUNE) will be exposed to a neutrino flux peaking at energies of a few GeV [3]. At these energies, one of the dominant neutrino interaction modes leads to the excitation of baryon resonances that subsequently decay producing pions. This process has never previously been measured for interactions.
This work presents the first measurement of the flux-averaged and charged-current (CC) single charged pion production cross section on argon using MicroBooNE. The final-state topology considered consists of an electron (or positron), a single charged pion, zero neutral pions (or heavier mesons), and any number of protons or neutrons:
@@
MicroBooNE collected data from two neutrino beams: the on-axis 8 GeV Booster Neutrino Beam and the approximately off-axis 120 GeV Neutrinos at the Main Injector (NuMI) beam [13]. In this work, data from the NuMI beam operating in a mixture of forward-horn-current (FHC) neutrino mode and reverse-horn-current (RHC) anti-neutrino mode are used. The NuMI flux at MicroBooNE contains a significant fraction of (%) due to the high beam energy and the significantly off-axis position of the detector. This makes it particularly useful for cross-section measurements. The average energies incident on MicroBooNE are MeV in FHC mode and MeV in RHC mode. The integrated exposure is protons-on-target (POT) in FHC mode and POT in RHC mode. This corresponds to the full NuMI dataset accumulated by MicroBooNE and used for the first time in this analysis.
@@
Simulation and reconstruction.— The NuMI neutrino flux is simulated with GEANT4 v4.10.4 [19, 20, 21, 22] constrained with available hadron production data using the PPFX software package [23]. The integrated flux of with energy over 60 MeV is /cm2/POT in FHC mode and /cm2/POT in RHC mode. The combined integrated flux is calculated by weighting the contributions from each horn-current mode according to the accumulated POT in that mode.
The LArSoft software framework [24] is used to perform simulation and reconstruction. Neutrino interactions are modeled using the GENIE v3.0.6 G18_10a_02_11a event generator [25] with the MicroBooNE tune applied [26]. In particular, resonant pion production is simulated using the Kuzmin-Lyubushkin-Naumov Berger-Sehgal model [27, 28, 29, 30] and coherent pion production using the Berger-Sehgal model [31]. Propagation of the final state particles through the detector is then performed using GEANT4 [19, 20, 21]. This is followed by simulation of the produced ionization electrons and scintillation light and the subsequent detector response [32, 33]. Simulated neutrino interactions are overlaid onto data collected while the beam is off providing data-based modeling of cosmic-ray induced interactions and detector noise. The Monte-Carlo (MC) prediction consists of simulated neutrino interactions within the detector, simulated interactions upstream of the detector (out-of-cryostat), and data collected with the beam off to model beam spills where no neutrino interaction occurs (EXT).
Reconstruction is performed using the Pandora pattern-recognition toolkit [34]. This uses a multi-algorithm approach to identify neutrino interactions from cosmic-ray-induced backgrounds and to reconstruct each resulting particle. Each particle associated with the neutrino interaction is categorized as a track (muons, pions, protons) or an electromagnetic shower (electrons, photons). Particle identification based on calorimetric and topological information is then performed [35, 36]. Finally, energy reconstruction is performed using particle range for tracks and through calorimetry for showers.
@@
The event selection expands upon tools developed in previous MicroBooNE analyses [9, 10, 37, 11]. Charged current interactions are first identified through the presence of an electromagnetic shower consistent with an electron. Tracks present in the event are then assumed to have originated from either protons or charged pions. These are distinguished based on their differing ionization profiles or the presence of pion re-interactions. The focus on first identifying charged current interactions avoids the need to distinguish between final state muons and charged pions that are otherwise more challenging to separate.
@@
Figure 2 shows the distribution of scores for the -rejection BDT and CC N-rejection BDT compared with data. The CC N-rejection BDT is shown after the cut on the -rejection BDT has been applied. Both are shown for FHC mode; RHC mode can be found in the Supplemental Material [40]. Both BDTs achieve good signal-background separation for their target topologies. In addition, excellent data–MC agreement within statistical and systematic uncertainties is seen across the full BDT score distributions.
Candidate CC interactions are selected with efficiencies of 6.2% and 5.8% and purities of 57.3% and 58.5% for the FHC and RHC beam periods, respectively. In total, 116 candidate events pass the selection in the combined POT data sample. The selected signal events are predicted to be 78% interactions and 22% interactions. They are dominated by resonant production (85%), with smaller contributions from deep-inelastic scattering (11%), quasi-elastic interactions (2%) and coherent interactions (2%).
@@
Systematic uncertainties are considered on: the neutrino flux from hadron production and beam-line geometry modeling [23, 22]; the neutrino interaction cross section modeling with GENIE [26]; secondary particle re-interactions [43]; detector response modeling including the scintillation light yield, recombination model, space charge effects [44], and ionization signal response [45]; out-of-cryostat interaction modeling; the number of argon targets; and POT counting. The dominant systematic uncertainty arises from the neutrino flux modeling (20–30%) due to the challenges in simulating the off-axis NuMI beam. This is followed by detector response modeling (15%), predominantly from recombination modeling due to the reliance on calorimetric variables to identify charged pions; and cross section modeling (10%). The other sources of systematic uncertainty are subdominant.
@@
The extracted differential cross sections in , , and are all consistent with the generator predictions. The preference for higher cross sections results in lower for NEUT, NuWro and GiBUU in , and ; whereas these models have a higher for hinting at possible shape disagreement in this variable. The largest are seen for driven by the smallest opening angle bin where, in particular, GENIE and GiBUU underpredict the cross section. However, due to the large statistical and flux uncertainties, all models lie within or close to 1 of the data suggesting in general good modeling of this process within the sensitivity of this measurement. To achieve greater model separation, future measurements would require greater statistics, potentially through improved reconstruction, and improved flux modeling. In particular, additional hadron production data to constrain the off-axis NuMI flux could significantly reduce the associated uncertainties.
Conclusions.— We have presented the first measurement of the flux-averaged charged current single charged pion production cross section on argon using the MicroBooNE detector. The full NuMI beam data set accumulated by MicroBooNE is used for the first time corresponding to a total exposure of POT with mean energy of 730 MeV. The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon. Differential cross sections are also reported as functions of electron energy, the electron and pion angles with respect to the neutrino direction, and the electron-pion opening angle. These are found to be in good agreement with generator predictions within uncertainties. This is the first measurement of pion production in interactions, one of the dominant interaction modes at the energies of the DUNE neutrino flux. It is essential to understand this process in order to facilitate the DUNE physics program, and direct measurements will have a critical impact to motivate modeling improvements. While this is the first time this process has been measured, the measurement is limited by statistical uncertainties and the challenges arising from the off-axis flux; future experiments in the SBN program [2] and the DUNE near detector [53] are expected to be able to improve on this.

### v2

@@ Abstract
Understanding electron neutrino interactions is crucial for measurements of neutrino oscillations and searches for new physics in neutrino experiments. We present the first measurement of the flux-averaged + charged current single charged pion production cross section on argon using the MicroBooNE detector and data from the NuMI neutrino beam. The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon at a mean + energy of 730 MeV. Differential cross sections are also reported in electron energy, electron and pion angles, and electron-pion opening angle.
Introduction.— The next generation of accelerator neutrino oscillation experiments will seek to address multiple open questions in neutrino physics through precision measurement of electron neutrino appearance in muon neutrino beams. These include the presence and scale of charge-parity violation in the neutrino sector, the neutrino mass ordering, and the resolution of anomalies observed at short baselines [1]. Several of these experiments will make use of the liquid argon time projection chamber (LArTPC) detector technology [2, 3]. As such, in order to facilitate these measurements, precise understanding of interactions on argon is essential.
Cross section modeling of interactions is typically extrapolated from measurements. However, uncertainties on the interaction cross section ratio arising from the different lepton masses and subsequent radiative corrections limit their constraining power [4, 5, 6, 7]. Understanding these effects through measurements of interactions is crucial for precision neutrino oscillation measurements, rare-event searches, and benchmarking theoretical models used in neutrino experiments.
@@
Existing measurements of cross sections on argon are limited and consist of several inclusive measurements [8, 9, 10] and a measurement without final state pions [11]. The Deep Underground Neutrino Experiment (DUNE) will be exposed to a neutrino flux peaking at energies of a few GeV [3]. At these energies, one of the dominant neutrino interaction modes leads to the excitation of baryon resonances that subsequently decay producing pions. This process has never previously been measured for interactions in argon.
This work presents the first measurement of the flux-averaged + charged-current (CC) single charged pion production cross section on argon using MicroBooNE. The final-state topology considered consists of an electron (or positron), a single charged pion, zero neutral pions (or heavier mesons), and any number of protons or neutrons:
@@
MicroBooNE collected data from two neutrino beams: the on-axis 8 GeV Booster Neutrino Beam and the approximately off-axis 120 GeV Neutrinos at the Main Injector (NuMI) beam [13]. In this work, data from the NuMI beam operating in a mixture of forward-horn-current (FHC) neutrino mode and reverse-horn-current (RHC) anti-neutrino mode are used. The NuMI flux at MicroBooNE is shown for each horn-current mode in the Supplemental Material [14]. The flux contains a significant fraction of (%) due to the high beam energy and the significantly off-axis position of the detector. This makes it particularly useful for cross-section measurements. The average energies incident on MicroBooNE are MeV in FHC mode and MeV in RHC mode. The integrated exposure is protons-on-target (POT) in FHC mode and POT in RHC mode. This corresponds to the full NuMI dataset accumulated by MicroBooNE and used for the first time in this analysis.
@@
Simulation and reconstruction.— The NuMI neutrino flux is simulated with GEANT4 v4.10.4 [20, 21, 22, 23] constrained with available hadron production data using the PPFX software package [24]. The integrated flux of with energy over 60 MeV is /cm2/POT in FHC mode and /cm2/POT in RHC mode. The combined FHC and RHC integrated flux is calculated by weighting the contributions from each horn-current mode according to the accumulated POT in that mode. The resulting combined integrated flux is /cm2/POT.
The LArSoft software framework [25] is used to perform simulation and reconstruction. Neutrino interactions are modeled using the GENIE v3.0.6 G18_10a_02_11a event generator [26] with the MicroBooNE tune applied [27]. In particular, resonant pion production is simulated using the Kuzmin-Lyubushkin-Naumov Berger-Sehgal model [28, 29, 30, 31] and coherent pion production using the Berger-Sehgal model [32]. Propagation of the final state particles through the detector is then performed using GEANT4 v4_10_3_p03c with the QGSP_BERT physics list [20, 21, 22]. This is followed by simulation of the produced ionization electrons and scintillation light and the subsequent detector response [33, 34]. Simulated neutrino interactions are overlaid onto data collected while the beam is off providing data-based modeling of cosmic-ray induced interactions and detector noise. The Monte-Carlo (MC) prediction consists of simulated neutrino interactions within the detector, simulated interactions upstream of the detector (out-of-cryostat), and data collected with the beam off to model beam spills where no neutrino interaction occurs (EXT).
Reconstruction is performed using the Pandora pattern-recognition toolkit. This uses a multi-algorithm approach to identify neutrino interactions from cosmic-ray-induced backgrounds and to reconstruct each resulting particle. Each particle associated with the neutrino interaction is categorized as a track (muons, pions, protons) or an electromagnetic shower (electrons, photons). A detailed description of the performance of the Pandora reconstruction in MicroBooNE can be found in Ref. [35]. Following the Pandora reconstruction, particle identification based on calorimetric and topological information [36, 37] is performed. Finally, energy reconstruction is performed using particle range for tracks and through calorimetry for showers.
@@
The event selection expands upon tools developed in previous MicroBooNE analyses [9, 10, 38, 11, 39]. Charged current interactions are first identified through the presence of an electromagnetic shower consistent with an electron. Tracks present in the event are then assumed to have originated from either protons or charged pions. These are distinguished based on their differing ionization profiles or the presence of pion re-interactions. The focus on first identifying charged current interactions avoids the need to distinguish between final state muons and charged pions that are otherwise more challenging to separate.
@@
Figure 2 shows the distribution of scores for the -rejection BDT and CC N-rejection BDT compared with data. The CC N-rejection BDT is shown after the cut on the -rejection BDT has been applied. Both are shown for FHC mode; RHC mode can be found in the Supplemental Material [14]. Both BDTs achieve good signal-background separation for their target topologies. For the rejection, 99.2% of induced photon showers are removed while keeping 42.0% of induced electron showers. For the CC N rejection, 94.3% of CC N interactions are rejected while keeping 67.1% of CC interactions. In addition, excellent data–MC agreement within statistical and systematic uncertainties is seen across the full BDT score distributions. A full description of the BDT input variables along with data–MC comparisons used can be found in the Supplemental Material [14].
Candidate CC interactions are selected with efficiencies of 6.2% and 5.8% and purities of 57.3% and 58.5% for the FHC and RHC beam periods, respectively. Further details on the selection performance can be found in the Supplemental Material [14]. In total, 116 candidate events pass the selection in the combined POT data sample. The selected signal events are predicted to be 78% interactions and 22% interactions. They are dominated by resonant production (85%), with smaller contributions from deep-inelastic scattering (11%), quasi-elastic interactions (2%) and coherent interactions (2%).
@@
Systematic uncertainties are considered on: the neutrino flux from hadron production and beam-line geometry modeling [24, 23]; the neutrino interaction cross section modeling with GENIE [27]; secondary particle re-interactions [44]; detector response modeling including the scintillation light yield, recombination model, space charge effects [45], and ionization signal response [46]; out-of-cryostat interaction modeling; the number of argon targets; and POT counting. In the unfolded results, the dominant systematic uncertainty arises from the neutrino flux modeling (20–30%) due to the challenges in simulating the off-axis NuMI beam. This is followed by detector response modeling (15%), predominantly from recombination modeling due to the reliance on calorimetric variables to identify charged pions; and cross section modeling (10%). The other sources of systematic uncertainty are subdominant.
@@
The extracted differential cross sections in , , and are all consistent with the generator predictions. The preference for higher cross sections results in lower for NEUT, NuWro and GiBUU in , and ; whereas these models have a higher for hinting at possible shape disagreement in this variable. The smallest are seen for where there is minimal shape difference between the data and the generator predictions. The largest are seen for driven by the smallest opening angle bin where, in particular, GENIE and GiBUU underpredict the cross section. However, due to the large statistical and flux uncertainties, all models lie within or close to 1 of the data suggesting in general good modeling of this process within the sensitivity of this measurement. To achieve greater model separation, future measurements would require greater statistics, potentially through improved reconstruction, and improved flux modeling. In particular, additional hadron production data to constrain the off-axis NuMI flux could significantly reduce the associated uncertainties.
Conclusions.— We have presented the first measurement of the flux-averaged charged current single charged pion production cross section on argon using the MicroBooNE detector. The full NuMI beam data set accumulated by MicroBooNE is used for the first time corresponding to a total exposure of POT with mean energy of 730 MeV. The total cross section is measured to be (0.93 0.13 (stat.) 0.27 (syst.)) cm2/nucleon. Differential cross sections are also reported as functions of electron energy, the electron and pion angles with respect to the neutrino direction, and the electron-pion opening angle. These are found to be in good agreement with generator predictions within uncertainties. This is the first measurement of pion production in interactions on argon, one of the dominant interaction modes at the energies of the DUNE neutrino flux. It is essential to understand this process in order to facilitate the DUNE physics program, and direct measurements will have a critical impact to motivate modeling improvements. While this is the first time this process has been measured, the measurement is limited by statistical uncertainties and the challenges arising from the off-axis flux; future experiments in the SBN program [2] and the DUNE near detector [54] are expected to be able to improve on this.